OpenStudy (anonymous):

what is the phase shift of y= 3cos(6x+ pi/2)

OpenStudy (anonymous):

pi/2

OpenStudy (anonymous):

be careful here

OpenStudy (anonymous):

if you want to read the phase shift from the formula it needs to look like \[A\cos(B(x-c))\] not \[A\cos(Bx -c)\]

OpenStudy (anonymous):

if you really want to graph this thing it is the same as \[-3\sin(6x)\] which it a lot easier to figure out. but to find the phase shift you will need to write it as \[3\cos(6(x+\frac{\pi}{12})\]

OpenStudy (anonymous):

so it would be \[\Pi/12\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

but i repeat it would be much easier to graph \[y=-3\sin(6x)\] where the amplitude is 3 and the period is \[\frac{\pi}{3}\]

OpenStudy (anonymous):

then to try to figure out what this looks like from the period and the phase shift.